OpenStudy (anonymous):

Find the derivative of xe^-2x

OpenStudy (anonymous):

@rajathsbhat can u help?

OpenStudy (anonymous):

can you do it if i teach you how to derive e^-2x?

OpenStudy (anonymous):

I got 2e^-2x is that correct?

OpenStudy (anonymous):

product rule..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what you got when you differentiated the exponent is correct......

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know the product rule??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this is the product rule..... it says if you have eg.... \[y=uv\] \[\frac{ dy }{ dx }= uv \prime + vu \prime \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you give an example?

OpenStudy (anonymous):

yes...sure eg. \[y= 7xe ^{x ^{2}}\] \[\frac{ dy }{ dx }=7x \times 2xe ^{x ^{2}} +xe ^{x ^{2}} \times 7\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you simplify it further from there?????

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

thx for the help

OpenStudy (anonymous):

you are welcome.....